OpenStudy (anonymous):

What is the value of the expression when a = 6 and b = 4? 14a-2[3b+] some one please help

OpenStudy (zaraxx):

i can't rlly do anything until my own question is answered so i can help u again if u like?

OpenStudy (anonymous):

oh sorry i dont need help right now but ill let u now cause im following the steps that u showed me

OpenStudy (zaraxx):

okie dokie! once u learn to input them and solve u've got it down :)

OpenStudy (anonymous):

can we go over it please

OpenStudy (anonymous):

@zaraxx

OpenStudy (zaraxx):

sure no problem. so we're going to input the values a=6 and b=4 into 14a-2(3b) which becomes 14*6-2(3*4)

OpenStudy (anonymous):

let me say what i got

OpenStudy (anonymous):

btw i will take this down when were done

OpenStudy (anonymous):

14x6=84

OpenStudy (anonymous):

right

OpenStudy (zaraxx):

yes, now what's next?

OpenStudy (anonymous):

84-2=82 right

OpenStudy (zaraxx):

yes, now start on the other side of the question.

OpenStudy (anonymous):

so 3x4 =12 and 12 +4 = 16 right

OpenStudy (zaraxx):

can u tell me what the original equation is again? is it (3b+b)?

OpenStudy (anonymous):

{14a-2[3b+4] a=6 b=4 } 14xa a=6 so 14 x6 =84 84-2=82 3xb b=4 so 3x4=12 12 +4=16 then what im stuck @zaraxx

OpenStudy (zaraxx):

well if it is 3b+b ur correct so far. now multiply 82 * 16 and that's ur answer

OpenStudy (anonymous):

82 x 16 =1312 right

OpenStudy (anonymous):

@ZARAXX

OpenStudy (zaraxx):

so since we had a bit of confusion im gonna go ahead and post the entire way to work it out so u know exactly how to do it later on. 14*6-2(3*4+4) 84-2*(16) 82 * 16 1312

OpenStudy (zaraxx):

exactly!! great job!!